# Determinantal random subgraphs

### v3

@@ 1.2. Further determinantal spanning subgraphs
In this paper, we define determinantal probability measures on spanning subgraphs with more complicated topology than trees (and cycle-rooted spanning forests), and give an explicit geometric formula for the weight of each graph appearing in the associated partition function. This is the content of Theorems 8.1 and 8.3. These random subgraphs are respectively spanning forests with fixed number of components and connected spanning subgraphs with fixed number of independent cycles, see Figure 1.111We note that all figures presented in this paper have been obtained using the well-known algorithm for sampling determinantal point processes (see [HKPV06] or [KL22, Proposition 3.13]). Unless otherwise specified, we use uniformly random chosen parameters for the samples in this paper.
@@
Defining a determinantal random subgraph is equivalent to specifying a self-adjoint positive contraction on the space of $1$-forms of the graph. The families discussed above correspond to the orthogonal projections on subspaces contained in, or containing, the space of exact forms (see Section 8). In the acyclic case (corresponding to subspaces contained in the space of exact forms), we further consider a certain operator canonically associated to the graph (a normalized Laplacian on $1$-forms) which is a convex combination of these (see Section 11.3) and we call the random spanning forest it induces, the Laplacian spanning forest (see Theorem 11.11).
@@ 1.4.1. Global-to-local method: perturbation of the partition function
    We define an operator $d_{\vartheta}=d+\omega_{\vartheta}$ whose image is $H=H_{0}\oplus{\mathrm{Vect}}(\theta_{1},\ldots,\theta_{k})$. By writing the squared volume of the image of an orthonormal basis under $d_{\vartheta}$ as the product of two terms (using Lemma 5.7), one of which involves only $d$, we obtain the ratio of this volume to the reference volume associated with $d$ as the squared norm of the orthogonal projection of  $\vartheta$ on $(H_{0}^{\perp})^{\wedge k}$ This proves Proposition 8.2.
@@ 1.4.2. Local-to-global method: perturbation of configuration weights
In the graphical case, simplifications occur due to the fact that invertible integer matrices have determinant equal to ±$\pm 1$. Without these simplifications in the general matroidal case, we need to carefully take care of new determinantal quantities.
@@ 1.5. Toolbox
Our main tools for proving these results are an exterior algebra version of the matrix-tree theorem (Propositions 3.13.3, and 13.1), and consequences of it, and the mean projection theorem for determinantal point processes (Theorem 5.5). The latter theorem was proven in [KL22, Theorem 5.9] and another proof was given in [KL23]; earlier instances of special cases of this statement appeared in [NS61], [Mau76, Theorem 1], [Big97, Proposition 7.3], [Lyo03, Proposition 6.8], and [CCK13, CCK15, Theorem A].
@@ 2.2.1. Chains and cycles
0 $0 \longrightarrow Z_{1}(\mathsf{G})\longrightarrow C_{1}(\mathsf{G})\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{\partial}}C_{0}(\mathsf{G})\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{\varepsilon}}\mathbb{Z}\longrightarrow 0$
@@ 2.2.2. Cochains and coboundaries
$0\longrightarrow\mathbb{Z}\longrightarrow C^{0}(\mathsf{G}) \mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{\delta}}C^{1}(\mathsf{G})\longrightarrow C^{1}(\mathsf{G})/B^{1}(\mathsf{G})\longrightarrow 0$
@@ 2.2.3. Subgraphs
Let us discuss cycles and coboundaries on subgraphs of $\mathsf{G}$, keeping in mind that, in this paper, all graphs are spanning, in the sense that their vertex set is ${\sf V}$. Thus, let us pick a subgraph $(S,{\sf V})$ of $\mathsf{G}$, that we simply refer to as $S$. There is no need for a special discussion of chains and cochains of degree 00 on $S$, since they are exactly the same as on $\mathsf{G}$. In degree $1$ however, the group of chains on $S$ identifies naturally with the submodule $C_{1}(S)\subseteq C_{1}(\mathsf{G})$ generated by the edges of $S$, and the group of cochains on $S$ with the quotient $C^{1}(\mathsf{G})/C_{1}(S)^{\circ}$, which in turn is isomorphic with the submodule $C^{1}(S)\subseteq C^{1}(\mathsf{G})$ generated by $\{e^{\star}:e\in S\}$.
@@
by close analogy with (10). The submodule $B^{1}(\mathsf{G},S)$ of $B^{1}(\mathsf{G})$ is the set of coboundaries of 00-chains that are locally constant on $S$, or more simply the set of coboundaries on $\mathsf{G}$ that vanish on $S$. In the classical terminology of algebraic topology, it would be called the group of relative coboundaries of the pair $(\mathsf{G},S)$. It will play for us the role of the group of coboundaries on the graph $\mathsf{G}_{/S}$ obtained by contracting all the edges of $S$, but we prefer not to consider this quotient graph and to keep working on $\mathsf{G}$.
@@ 2.3.1. Chains on $\mathsf{G}$
To give a more concrete expression of ${\sf Z}_{T}$, given two vertices $v,w$ of $\mathsf{G}$, let us denote by $[v,w]_{T}$ the unique simple path going from $v$ to $w$ in $T$, seen as an element of $C_{1}(\mathsf{G})$. This is the unique element of $C_{1}(T)$ such that $\partial([v,w]_{T})=w-v$. Then for each edge $e$, we have $\partial e=-\partial([{\sf{t}}(e),{\sf{s}}(e)]_{T})$, so that the abovementioned splitting reads, in this case, $e=-[{\sf{t}}(e),{\sf{s}}(e)]_{T}+(e+[{\sf{t}}(e),{\sf{s}}(e)]_{T})$. Therefore,
@@ 3. Integral multilinear identities
We denote the exterior algebra of a $\mathbb{Z}$-module $M$ by $M$. In $\,C_{1}(\mathsf{G})$ and $\,C^{1}(\mathsf{G})$, we will use the following notation : for every subset $S$ of ${\sf E}$, with $S\cap{\sf E}^{+}=\{e_{1},\ldots,e_{r}\}$ enumerated in the order of ${\sf E}^{+}$, we write
(24) $e_{S}=e_{1}\wedge\ldots\wedge e_{r}\ \text{ and }\ e^{\star}_{S}=e^{\star}_{1}\wedge\ldots\wedge e^{\star}_{r}.$
@@ 3.1. Connected spanning subgraphs
Let us choose an integer $k\geqslant 0$ and a subgraph $K\in\mathcal{C}_{k}(\mathsf{G})$ of $\mathsf{G}$, that is, a connected spanning subgraph of $\mathsf{G}$ with $k$ linearly independent cycles, in the sense that the $\mathbb{Z}$-module $Z_{1}(K)$ has rank $k$. Let us fix an arbitrary integral basis $\mathscr{Z}^{K}$ of $Z_{1}(K)$. Our two main identities give an expression of the element $\det\mathscr{Z}^{K}$ of $^{k}C_{1}(\mathsf{G})$, and an expression of its tensor square.
@@
In $^{k}C_{1}(\mathsf{G})$, the following identity holds:
@@
Let us write $\mathscr{Z}^{K}=(z_{1},\ldots,z_{k})$ and decompose the element $z_{1}\wedge\ldots\wedge z_{k}$ of $^{k}C_{1}(\mathsf{G})$ on the basis $\{e_{S}:|S|=k\}$:
@@
Moreover, since $^{k}Z_{1}(K)$ is a module of rank $1$ generated by $\det\mathscr{Z}^{K}$, the last equation combined with (25) implies that on the submodule $^{k}Z_{1}(K)$ of $^{k}C_{1}(\mathsf{G})$,
@@
In ($^{k}C_{1}(\mathsf{G})\big{)}^{\otimes 2}$, the following equality holds:
@@ 3.2. Acyclic spanning subgraphs
In $^{k}C^{1}(\mathsf{G})$, the following identity holds:
@@
Then, since $^{k}B^{1}(\mathsf{G},F)$ is a module of rank $1$ generated by $\det\mathscr{B}^{F}$, the last identity combined with (28) implies that on the submodule $^{k}B^{1}(\mathsf{G},F)$ of $^{k}C^{1}(\mathsf{G})$,
@@
In ($^{k}C^{1}(\mathsf{G})\big{)}^{\otimes 2}$, one has
@@ 4. Metric structures on cochains
$\Omega^{i}(\,\cdot\,)=C^{i}(\,\cdot\,)\otimes\mathbb{K}\ \text{ and }\ \Omega_{i}(\,\cdot\,)=C_{i}(\,\cdot\,)\otimes\mathbb{K}.$
We will use cochains with coefficients in $\mathbb{K}$ more often than chains, and instead of cochains we will often call them forms or, to be more precise, 00-forms and $1$-forms.
@@ 4.3. Exterior powers
For every $k\geqslant 0$, the family $\{e^{\star}_{S}:S\subseteq{\sf E}^{+},|S|=k\}$ is an orthogonal basis of $^{k}\Omega^{1}(\mathsf{G})$, with
@@ 4.4. Inner products on chains
We chose not to endow $\Omega_{0}(\mathsf{G})$ nor $\Omega_{1}(\mathsf{G})$ with an inner product. If we had to, we would take the one which turns the maps ${{\sf J}^{1}_{\underline{x}}}$ and ${\sf J}^{0}$ into isometries. On $\Omega_{0}(\mathsf{G})$, this would be the canonical inner product, since there is no dependency on $\underline{x}$ there. On $\Omega_{1}(\mathsf{G})$, this would be the inner product for which the canonical basis is orthogonal and for which $\|e\|^{2}=x_{e}^{-1}$ for every edge $e$. Under this inner product, for every $k\geqslant 0$, the family $\{e_{S}:S\subseteq{\sf E}^{+},|S|=k\}$ is orthogonal in $^{k}\Omega_{1}(\mathsf{G})$, and for each $S\subseteq{\sf E}^{+}$,
@@ 4.7. Real tori and finite abelian groups
Comparing with the previous equality, we deduce that all signs in the last sum are the same.444 The fact that all signs are the same in the above formula is also a reflection of the fact that ${{\sf J}^{1}_{\underline{x}}}^{\wedge b_{1}}\det\mathscr{Z}$ and $\det\mathscr{B}$ are Hodge dual of each other, up to a constant. Thus, we proved the following proposition.
@@
Let $\mathscr{Z}$ be a basis of $Z_{1}(\mathsf{G})$ and $\mathscr{B}$ a basis of $B^{1}(\mathsf{G})$. Then in the line $^{|{\sf E}^{+}|}\Omega^{1}(\mathsf{G})$, we have

### v4

@@ 1.2. Further determinantal spanning subgraphs
In this paper, we define determinantal probability measures on spanning subgraphs with more complicated topology than trees (and cycle-rooted spanning forests), and give an explicit geometric formula for the weight of each graph appearing in the associated partition function. This is the content of Theorems 8.1 and 8.3. These random subgraphs are respectively spanning forests with fixed number of components and connected spanning subgraphs with fixed number of independent cycles, see Figure 1.111We note that all figures presented in this paper have been obtained using the well-known algorithm for sampling determinantal point processes (see [HKPV06] or [KL22, Proposition 3.13]). Unless otherwise specified, we use uniformly random chosen parameters for the samples in this paper. Our basic sampling code is available on the repository https://plmlab.math.cnrs.fr/thierrylevy/determinantal-random-subgraphs. As pointed out to us by one of the referees, it is possible to use faster sampling strategies since the determinantal point processes considered here have fixed cardinality, such as the one of [BTA23]. Letting $e$ and $v$ be the number of edges and vertices, this would yield a complexity of $O(e(v-1-k)^{2})$ rather than $O(e^{3})$ which in our examples would not change much since $e$ is of the order of $v$ and $k$ is small. Moreover, we are unaware of a Wilson style algorithm [Wil96] for the probability measures we consider, but having one would naturally be of great use. As will be apparent from (1) and (4), let us note right away that these probability measures are not ‘uniform’ [^2] Uniform in the sense of being proportional to the product of edge weights of each configuration. unless we are considering the intersection of these families, namely the case of spanning trees.
@@
Defining a determinantal random subgraph is equivalent to specifying a self-adjoint positive contraction on the space of $1$-forms of the graph. The families discussed above correspond to the orthogonal projections on subspaces contained in, or containing, the space of exact forms (see Section 8). [^3] The determinantal random spanning tree corresponds to the space where this subspace is the space of exact $1$-forms. Surprisingly, it also is the ‘uniform’ measure, by the fundamental result of Burton and Pemantle. In that sense, these families generalize the ‘uniform’ spanning tree case, although, as already said, the other measures we consider in this paper are generically not ‘uniform’. In the acyclic case (corresponding to subspaces contained in the space of exact forms), we further consider a certain operator canonically associated to the graph (a normalized Laplacian on $1$-forms) which is a convex combination of these (see Section 11.3) and we call the random spanning forest it induces, the Laplacian spanning forest (see Theorem 11.11).
@@ 1.4.1. Global-to-local method: perturbation of the partition function
    We define an operator $d_{\vartheta}=d+\omega_{\vartheta}$ whose image is $H=H_{0}\oplus{\mathrm{Vect}}(\theta_{1},\ldots,\theta_{k})$. By writing the squared volume of the image of an orthonormal basis under $d_{\vartheta}$ as the product of two terms (using Lemma 5.7), one of which involves only $d$, we obtain the ratio of this volume to the reference volume associated with $d$ as the squared norm of the orthogonal projection of  $\vartheta$ on $(H_{0}^{\perp})^{\wedge k}$ . This proves Proposition 8.2.
@@ 1.4.2. Local-to-global method: perturbation of configuration weights
In the graphical case, simplifications occur due to the fact that invertible integer matrices have determinant equal to $\pm 1$. Without these simplifications in the general matroidal case, we need to carefully take care of new determinantal quantities.
@@ 1.5. Toolbox
Our main tools for proving these results are an exterior algebra version of the matrix-tree theorem (Propositions 3.13.3, and 13.1), and consequences of it, and the mean projection theorem for determinantal point processes (Theorem 5.5). The latter theorem was proven in [KL22, Theorem 5.9] and another proof was given in [KL23]; earlier instances of special cases of this statement appeared in [NS61], [Mau76, Theorem 1], [Big97, Proposition 7.3], [Lyo03, Proposition 6.8], and [CCK13, CCK15, Theorem A]. As pointed out to us by one of the referees, a form of this theorem also appears in [DM21, Theorems 2 and 3].
@@ 2.2.1. Chains and cycles
[…]
@@ 2.2.2. Cochains and coboundaries
$0\longrightarrow\mathbb{Z}\longrightarrow C^{0}(\mathsf{G}) \mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{\delta}}C^{1}(\mathsf{G})\longrightarrow C^{1}(\mathsf{G})/B^{1}(\mathsf{G})\longrightarrow 0.$
@@ 2.2.3. Subgraphs
Let us discuss cycles and coboundaries on subgraphs of $\mathsf{G}$, keeping in mind that, in this paper, all graphs are spanning, in the sense that their vertex set is $\sf V$}. Thus, let us pick a subgraph $(S,{\sf V})$ of $\mathsf{G}$, that we simply refer to as $S$. There is no need for a special discussion of chains and cochains of degree 0 on $S$, since they are exactly the same as on $\mathsf{G}$. In degree $1$ however, the group of chains on $S$ identifies naturally with the submodule $C_{1}(S)\subseteq C_{1}(\mathsf{G})$ generated by the edges of $S$, and the group of cochains on $S$ with the quotient $C^{1}(\mathsf{G})/C_{1}(S)^{\circ}$, which in turn is isomorphic with the submodule $C^{1}(S)\subseteq C^{1}(\mathsf{G})$ generated by $\{e^{\star}:e\in S\}$.
@@
by close analogy with (10). The submodule $B^{1}(\mathsf{G},S)$ of $B^{1}(\mathsf{G})$ is the set of coboundaries of 0-chains that are locally constant on $S$, or more simply the set of coboundaries on $\mathsf{G}$ that vanish on $S$. In the classical terminology of algebraic topology, it would be called the group of relative coboundaries of the pair $(\mathsf{G},S)$. It will play for us the role of the group of coboundaries on the graph $\mathsf{G}_{/S}$ obtained by contracting all the edges of $S$, but we prefer not to consider this quotient graph and to keep working on $\mathsf{G}$.
@@ 2.3.1. Chains on $\mathsf{G}$
To give a more concrete expression of ${\sf Z}_{T}$, given two vertices $v,w$ of $\mathsf{G}$, let us denote by $v,w]_{T}$ the unique simple path going from $v$ to $w$ in $T$, seen as an element of $C_{1}(\mathsf{G})$. This is the unique element of $C_{1}(T)$ such that $\partial([v,w]_{T})=w-v$. Then for each edge $e$, we have $\partial e=-\partial([{\sf{t}}(e),{\sf{s}}(e)]_{T})$, so that the abovementioned splitting reads, in this case, $e=-[{\sf{t}}(e),{\sf{s}}(e)]_{T}+(e+[{\sf{t}}(e),{\sf{s}}(e)]_{T})$. Therefore,
@@ 3. Integral multilinear identities
[…]
@@ 3.1. Connected spanning subgraphs
Let us choose an integer $k\geqslant 0$ and a subgraph $K\in\mathcal{C}_{k}(\mathsf{G})$ of $\mathsf{G}$, that is, a connected spanning subgraph of $\mathsf{G}$ with $k$ linearly independent cycles, in the sense that the $\mathbb{Z}$-module $Z_{1}(K)$ has rank $k$. Let us fix an arbitrary integral basis $\mathscr{Z}^{K}$ of $Z_{1}(K)$. Our two main identities give an expression of the element $\det\mathscr{Z}^{K}$ of $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}C_{1}(\mathsf{G})$, and an expression of its tensor square.
@@
In $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}C_{1}(\mathsf{G})$, the following identity holds:
@@
Let us write $\mathscr{Z}^{K}=(z_{1},\ldots,z_{k})$ and decompose the element $z_{1}\wedge\ldots\wedge z_{k}$ of $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}C_{1}(\mathsf{G})$ on the basis $\{e_{S}:|S|=k\}$:
@@
Moreover, since $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}Z_{1}(K)$ is a module of rank $1$ generated by $\det\mathscr{Z}^{K}$, the last equation combined with (25) implies that on the submodule $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}Z_{1}(K)$ of $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}C_{1}(\mathsf{G})$,
@@
In $\big(\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}C_{1}(\mathsf{G})\big)^{\otimes 2}$, the following equality holds:
@@ 3.2. Acyclic spanning subgraphs
In $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}C^{1}(\mathsf{G})$, the following identity holds:
@@
Then, since $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}B^{1}(\mathsf{G},F)$ is a module of rank $1$ generated by $\det\mathscr{B}^{F}$, the last identity combined with (28) implies that on the submodule $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}B^{1}(\mathsf{G},F)$ of $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}C^{1}(\mathsf{G})$,
@@
In $\big(\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}C^{1}(\mathsf{G})\big)^{\otimes 2}$, one has
@@ 4. Metric structures on cochains
[…]
We will use cochains with coefficients in $\mathbb{K}$ more often than chains, and instead of cochains we will often call them forms or, to be more precise, 0-forms and $1$-forms.
@@ 4.3. Exterior powers
For every $k\geqslant 0$, the family $\{e^{\star}_{S}:S\subseteq{\sf E}^{+},|S|=k\}$ is an orthogonal basis of $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}\Omega^{1}(\mathsf{G})$, with
@@ 4.4. Inner products on chains
We chose not to endow $\Omega_{0}(\mathsf{G})$ nor $\Omega_{1}(\mathsf{G})$ with an inner product. If we had to, we would take the one which turns the maps ${\sf J}^{1}_{\underline{x}}$} and ${\sf J}^{0}$ into isometries. On $\Omega_{0}(\mathsf{G})$, this would be the canonical inner product, since there is no dependency on $\underline{x}$ there. On $\Omega_{1}(\mathsf{G})$, this would be the inner product for which the canonical basis is orthogonal and for which $\|e\|^{2}=x_{e}^{-1}$ for every edge $e$. Under this inner product, for every $k\geqslant 0$, the family $\{e_{S}:S\subseteq{\sf E}^{+},|S|=k\}$ is orthogonal in $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}\Omega_{1}(\mathsf{G})$, and for each $S\subseteq{\sf E}^{+}$,
@@ 4.7. Real tori and finite abelian groups
Comparing with the previous equality, we deduce that all signs in the last sum are the same. [^6] The fact that all signs are the same in the above formula is also a reflection of the fact that ${{\sf J}^{1}_{\underline{x}}}^{\wedge b_{1}}\det\mathscr{Z}$ and $\det\mathscr{B}$ are Hodge dual of each other, up to a constant. Thus, we proved the following proposition.
@@
Let $\mathscr{Z}$ be a basis of $Z_{1}(\mathsf{G})$ and $\mathscr{B}$ a basis of $B^{1}(\mathsf{G})$. Then in the line $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{|{\sf E}^{+}|}\Omega^{1}(\mathsf{G})$, we have